Here are summary statistics for randomly selected weights of newborn​ girls: nequals211​, x overbarequals26.8 ​hg, sequals6.2 hg. Construct a confidence interval estimate of the mean. Use a 99​% confidence level. Are these results very different from the confidence interval 24.9 hgless thanmuless than27.9 hg with only 15 sample​ values, x overbarequals26.4 ​hg, and sequals1.9 ​hg? What is the confidence interval for the population mean mu​? nothing hgless thanmuless than nothing hg ​(Round to one decimal place as​ needed.) Are the results between the two confidence intervals very​ different? A. ​Yes, because one confidence interval does not contain the mean of the other confidence interval. B. ​No, because the confidence interval limits are similar. C. ​Yes, because the confidence interval limits are not similar. D. ​No, because each confidence interval contains the mean of the other confidence interval.
Answer #1

Given that,

Point estimate = sample mean = = 26.8 hg

sample standard deviation = s = 6.2 hg

sample size = n = 211

Degrees of freedom = df = n - 1 = 211 - 1 = 210

At 99% confidence level

= 1 - 99%

=1 - 0.99 =0.01

/2 = 0.005

t/2,df = t0.005, 210 = 2.599

Margin of error = E = t/2,df * (s /n)

= 2.599 * (6.2 / 211)

Margin of error = E = 1.1

The 99% confidence interval estimate of the population mean is,

- E < < + E

26.8 - 1.1 < < 26.8 + 1.1

(25.7 hg < < 27.9 hg)

B)   Yes, because the confidence interval limits are not similar